OpenStudy (anonymous):

C=5over9(F-32) is used to find temp. C is Celsius and F is Fahrenheit. use the equation to find the equivalent F temp for 70 degrees C the formula is: F=__, and 70 degrees C=__ degrees F i got 158 degrees for F but can not figure out C and I dont know if 158 is even right for F

OpenStudy (hoblos):

C=(5/9)(F-32) (9/5)C = F-32 then F = (9/5)C +32 for C = 70 F= (9/5)(70) +32 = 126+32 = 158

OpenStudy (anonymous):

so whats c haha math sucks!